y = ģ ( - 32+2) -. (2 f(3) = xen 5 i fix) 2 5xAnalyze and sketch the graph of each function. Label intercepts, relative extrema Points of Inflection and asym proles. work
Answer #1

Solution U Da y = (re?-32+2) 3 C-00,00) Domain the of gives function is then Now put y=0 find intercepts - cs put x=0 y = 2 t- M Range of 22-32+2 veutere. (xw, yu) it a so the range is f(2)= qu if a co range is f(x) = you the f(x) = -1 then the Rangefc2l= I ker Domain of fire)= + ker 15 10 0 54 2 2 -10 F-15 Houe 201 ou x>0 Domain of f(x) = (-0,0) 1 (0,00) critical point =Asymprotes of y = f(x) = 1 here veutical Asymptotes of f(x) zez xex_o X=0 Houizontal Asymprotes of fire) 4=0 Symmetry of f(x)3) (-00,00) cij =0. f(x)= x5_52 Is Domain of function find intercept points dif x=0 y=0 (0,0) if you 25-5x=0 x(ne-5)=0 +415.6 4 inflection points f(x)=75-6% plcx)= 524-5 f(x)= 20C3 inflection points are in Domain edges, gon8=o Heue mot on x=0 y=0 L